Grade 12Physical Chemistry

The values of Ksp of two sparingly soluble salts Ni(OH)2 and AgCN are 2.0 × 10–15 and 6 × 0–17 respectively. Which salt is more soluble? Explain ?

sudhanshu
12 Years agoGrade 12

1 Answer

Gaurav
11 Years ago
Hello Student

AgCN ⇌ Ag+ + CN
Ksp = [Ag+ ][CN ] = 6 × 10–17
Ni(OH)2 ⇌ Ni2+ + 2OH
Ksp = [Ni2+ ][OH– ] 2 = 2 × 10–15
Let [Ag+ ] = S1 , then [CN- ] = S1
Let [Ni2+ ] = S2 , then [OH– ] = 2S2
S12 = 6 × 10–17 ,
S1 = 7.8 × 10–9 (S2 )(2S2 )2 = 2 × 10–15,
S2 = 0.58 × 10–4
Ni(OH)2 is more soluble than AgCN